OpenStudy (anonymous):

How do I solve this y′′−3y′+2y=3e^(-x)-10cos(3x) y(0)=5, y'(0)=-2? i can solve the first part of the equation by using the homogeneous solutions of c1e^x+c2e^2x+yp=0 and but when i try solving the yp i get no were.

OpenStudy (loser66):

for non-homogenous part, you have 2terms, right?, so solve first term first It it 3e^(-x) . That is y'' -3y'+2y = 3e^(-x) the same with the second term, then combine them all.

OpenStudy (loser66):

for homogenous part, why don't you replace y(0) and y'(0) to find C1, C2?

OpenStudy (anonymous):

i thought i had to solve for a general soultion before i input my iv

OpenStudy (anonymous):

oh if i seperate them would i be doing the undetermined coefficients methods for both of them?

OpenStudy (loser66):

I think you misunderstand the method!! You have 2 parts: homogenous y"-3y'+2y =0 , that will give you 2 homogeneous solutions, let call y_g1 and y_g2 non homogenous part , that I showed you above, that will give you some other solutions. then. combine them all

OpenStudy (anonymous):

i think im lost

OpenStudy (loser66):

For homogenous part: Characteristic equation is r^2 -3r +2 = 0 , r = 2 and r =1 that gives you the solutions: y_g1 = C1 e ^2x and y_g2 = C2 e^x

OpenStudy (loser66):

Now, plug y(0) = 5 and y'(0) = -2 you have C1, C2, what are they?

OpenStudy (anonymous):

i got my homogeneous solution but for the second part the 3e^(-x)-10cos3x i cant solve

OpenStudy (anonymous):

c1=-7 and c2=12

OpenStudy (amorfide):

you use particular integrals for them

OpenStudy (loser66):

ok, let amorfide cooks this meal. :)

OpenStudy (anonymous):

ok. When you mean particular integrals do you mean when you set cx=ae^(-x)cos3x+be^(-x)sin3x and take two derivatives of it to get c'' and c'?

OpenStudy (anonymous):

and then pop those results into the original equation to solve

OpenStudy (amorfide):

Particular integral function for \[e ^{-x}\] is \[y=Ce ^{-x}\] but since you already have the form \[Ae ^{-x}\] in your complimentary function you must multiply our particular integral by a factor of x to get \[y=Cxe ^{-x}\] from here you will differentiate it twice then substitute y, first derivative, second derivative into your original second ODE you will then solve for C you will then substitute C into your particular integral to get Y for example, if C=3 you get \[y=3e ^{-x}\] this is added to your complimentary function next you have to do the same method with the particular integral for cos(x) particular integral for cos(3x) is \[y= C _{1}sinx + C_{2}cosx\] once you work out this particular integral, add this onto your complimentary function btw your complimentary function is \[y=Ae^{2x} + Be^{x}\] your general solution is \[y=Ae^{2x} + Be^{x}\] + particular integral for \[e^{-x}\] + particular integral for cos(3x)

OpenStudy (amorfide):

I feel like that is explained pretty poorly

OpenStudy (anonymous):

no i was going over what you said and im inputting the information to see if i can get it. thanks im going to try that out thanks.